### v1

# FlowMotion: Target-Predictive Flow Matching for Realistic Text-Driven Human Motion Generation

Abstract
Achieving highly diverse and perceptually consistent 3D character animations with natural motion and low computational costs remains a challenge in computer animation. Existing methods often struggle to provide the nuanced complexity of human movement, resulting in perceptual inconsistencies and motion artifacts. To tackle these issues, we introduce FlowMotion, a novel approach that leverages Conditional Flow Matching (CFM) for improved motion synthesis. FlowMotion incorporates an innovative training objective that more accurately predicts target motion, reducing the inherent jitter associated with CFM while enhancing stability, realism, and computational efficiency in generating animations. This direct prediction approach enhances the perceptual quality of animations by reducing erratic motion and aligning the training more closely with the dynamic characteristics of human movement. Our experimental results demonstrate that FlowMotion achieves higher balance between motion smoothness and generalization capability while maintaining the computational efficiency inherent in flow matching compared to state-of-the-art methods.
keywords:
3D Animation , Human Motion Synthesis, Flow Matching
1 Introduction
The synthesis of 3D human body motion has diverse applications across fields such as robotics [1, 2, 3], VR/AR [4, 5], entertainment [6, 7, 8, 9], and social interaction in virtual 3D spaces [10, 11, 12, 13]. To achieve realistic and context-aware motion synthesis, motion generation techniques frequently leverage data-driven motion capture data. These techniques include, for instance, approaches based on deep learning [14, 15], reinforcement learning [16, 17], and hybrid methods [18], which utilize these datasets to model and predict complex movements.
A significant subset of techniques for synthesizing 3D human body motion has emerged, driven by recent advances in generative models. These methods have efficiently enabled the generation of motion conditioned on user-defined inputs, with a primary focus on descriptive texts, that specifies the intended movement. Such approaches leverage the adaptability of generative models to produce contextually appropriate and plausible motion sequences  [19, 20, 21, 22, 23, 24]. While recent advances in 3D human motion generation are significant, several challenges remain. The inherent complexities of motion generation are exacerbated by the need to incorporate diverse constraints, such as spatial trajectories [25, 26], interactions with surrounding objects [27, 28], or temporal specifications defined by keyframes [29], all aimed at producing lifelike movements.
Among these generative methods, classes such as variational autoencoders (VAEs) [21, 22, 30], generative adversarial networks (GANs) [31, 23, 32, 24], and diffusion models [20, 19, 33] have received great attention for 3D human motion generation, each offering distinct advantages and limitations. While VAEs have been used in motion generation, their output diversity is often limited due to the issue of posterior collapse. Similarly, GANs can suffer from mode collapse, making it challenging to achieve sufficient diversity in the generated motions. In contrast, diffusion models excel in generating diverse and natural motions, albeit with increased latency during the generation process.
To address these limitations, hybrid models have emerged to leverage the strengths of different architectures. Recent proposals include integrating features from GANs and diffusion models to enhance generation speed without sacrificing generalization capability [34]. Furthermore, there are also approaches combining the strengths of GANs, VAEs, and diffusion models, utilizing distillation techniques for even faster sampling [35]. Additionally, diffusion techniques have been applied to latent spaces to enable faster sampling, using VAEs [36].
One of the main advantages of generative methods is their capability of producing a diverse range of plausible human motion sequences, enabling users to explore multiple interpretations of a desired motion and select the sequence that best aligns with their creative intent. This flexibility is particularly valuable in scenarios requiring user-controlled motion generation.
Beyond the aforementioned generative approaches, flow-matching-based models have demonstrated promising performance in generating both images [37, 38] and 3D human motion [39], achieving fast synthesis with adequate diversity in the results. However, in our analysis of the flow matching approach by Hu et al. [39] for 3D human body motion synthesis, we observed temporal inconsistencies (jitter effect) [4], which can compromise smoothness and perceptual fidelity. We hypothesize that this artifact arises from the conditional flow matching (CFM) formulation [37], where the calculation of the learned vector field inherently incorporates variations resulting from the differences between the data representations and samples from a standard Gaussian distribution used during training.
To take advantage of the computational efficiency of flow matching while avoiding the jitter effect, we propose a novel training objective explicitly designed to approximate the original, unperturbed motion, effectively attenuating fluctuations within the vector field during synthesis. Additionally, we employ the acceleration metric to rigorously quantify the temporal smoothness of the synthesized motion and ensure a high degree of fidelity in the generated results.
Our approach presents the following main contributions:
  • 1.
    Target-oriented training objective: We introduce a novel, target-oriented training objective for CFM that directly predicts the target motion. This approach outperforms diffusion and flow matching methods in temporal coherence and fidelity.
  • 2.
    Temporal smoothness and jitter reduction: By introducing a novel training objective and leveraging the acceleration metric, we ensure smooth transitions and minimize undesirable tremors in the generated motion.
  • 3.
    Balance between fidelity and smoothness: Our approach achieves a superior trade-off between generation fidelity, evaluated using Fréchet Inception Distance (FID), and temporal smoothness, assessed by jitter metrics (Fig. 1), resulting in reliable and natural motion sequences.
This paper is organized as follows: In Sec. 2, we review the related work on generative methods for 3D human motion synthesis. In Sec. 3, we present our method, covering both theoretical foundations and computational details. Sec. 4 describes the experiments conducted to evaluate the effectiveness of our approach. In Sec. 5, we discuss and analyze the results of these experiments. Finally, we conclude our study in Sec. 6, highlighting potential directions for future work based on the insights gained from our approach.
Figure 1: FID-Jitter evaluation on the HumanML3D dataset. FlowMotion achieves a favorable balance between motion smoothness, as quantified by lower Jitter, and generation fidelity, as measured by lower Fréchet Inception Distance (FID) scores, demonstrating its ability to generate high-quality motion compared to prior methods.
2 Related Work
We review techniques for 3D human motion generation in two subsections. Section 2.1 provides an overview of generative models, ranging from GANs and VAEs to recent Transformer-based architectures, discussing their challenges and corresponding contributions. Section 2.2 focuses on diffusion models, which have garnered significant attention and improvements, and flow matching models due to their shared characteristics.
2.1 3D Human Motion Generation
Diverse studies have explored the synthesis of human body movements conditioned on text [20, 30, 19]. For instance, GANs have been used for generating movements based on specific actions described in text [23]. Similarly, Text2Action [24] employs a Seq2Seq-inspired model [40] to translate text into motion, focusing on upper body movements.
More recently, Guo et al. [30] introduced a temporal VAE enhanced with attention mechanisms for generating human motion from text, alongside the widely adopted HumanML3D dataset and associated evaluation metrics. This dataset, one of the largest currently available, has become a standard benchmark for subsequent research, including the present work.
Subsequent advancements have seen the incorporation of Transformer architectures. ACTOR [41] utilized a Transformer-based VAE, while TEACH [42] conditioned motion generation on sequences of textual motion descriptions using a similar architectural approach. Further, TM2T [43] established connections between text tokens and motion sequences via Transformers, demonstrating efficacy on both HumanML3D and KIT-ML [44] datasets. T2M-GPT [21] employed a Vector Quantized Variational AutoEncoder (VQ-VAE) in conjunction with a CLIP text encoder [45] to achieve a discrete representation of movement, exhibiting robust generalization capabilities. Text2Gestures [46] utilizes a Transformer encoder to process descriptive text encoded with GloVe [47] and a Transformer decoder which, combined with previous movements, generates subsequent ones, enabling the production of expressive animations.
2.2 Diffusion and Flow Matching Models
Diffusion models, based on stochastic processes, have achieved significant success in tasks such as image [48, 49], sound [50, 51], and video generation [52]. Recently, several methods have leveraged diffusion models for motion generation [20, 19]. Among these, MDM [20] employs a Transformer encoder to generate motion sequences through iterative denoising. The model optimizes a loss function that directly targets the final, denoised motion state. MDM integrates CLIP for text encoding and has been evaluated on standard benchmarks, supporting both text-conditioned and unconditional generation. Its architecture has been widely adopted, including in work by Cohan et al. [26], which extends it to generate motions constrained by keyframes and trajectories.
Likewise, MotionDiffuse [19] employs a more extensive network than MDM, using a Transformer encoder for noisy motion input and a Transformer decoder, specifically a cross-attention mechanism, for handling the input text. It uses CLIP for text encoding and is evaluated on established datasets. A similar approach is taken in FLAME [53], which uses RoBERTa [54] for text encoding. Recently, attention networks combined with databases to retrieve movements that approximate textual input descriptions have been used alongside diffusion models to achieve better generalization in motion generation [55]. MoFusion [33], based on diffusion, does not use a Transformer for processing sequential motion data but instead employs a 1D-UNet, similar to that used in Stable Diffusion [56] for images. It extends the generation process to be conditioned not only on text but also on audio.
However, these diffusion-based approaches are often slow in the inference process due to the extensive sequence of steps required for sampling. Efforts like MLD [36] improve inference speed by performing diffusion in a latent space achieved through a prior VAE process, reducing computational resources and accelerating the inference process. Similarly, MotionLCM [35] and EMDM [34] achieve real-time inference by accelerating the sampling process, skipping denoising steps in the diffusion process through distillation and a discriminator process similar to those used in GANs, respectively. To ensure coherent body generation and prevent artifacts such as levitation, PhysDiff [18] integrate physical constraints during the sampling phase of the diffusion framework, though such constraints may reduce generation speed.
On the other hand, text-conditioned human motion generation via flow matching has shown promising results [39]. Using the same architecture as MDM, this approach achieves better generalization and faster inference due to the streamlined trajectory in the sampling process. However, the use of flow matching relies on intractable integrals. Thus, in practice, one uses the Conditional Flow Matching (CFM) model [37], but it tends to generate erratic or jittering motions during inference. The variations introduced by the CFM formulation, where the learned vector field is influenced by the interpolation between data representations and random Gaussian noise, may contribute to this artifact. In our method, we propose reducing jittering by modifying the loss function of the CFM, such that it directly compares the generated image with the noise-free original image.
Distinct from previous work, our results indicate that the technique here presented delivers visually superior results without increasing computational cost.
3 The FlowMotion Method
Our method generates realistic human motion sequences conditioned on textual descriptions, with focus on producing smooth and controlled movements. To achieve this, we leverage Conditional Flow Matching (CFM) [37], which enables the learning of complex motion distributions. In contrast to standard flow matching approaches, CFM operates on conditional distributions, which is advantageous for sampling as it promotes straighter trajectories and therefore, faster sampling.
Our training objective directly predicts the target motion, enhancing stability and reducing jitter compared to prior CFM and diffusion-based approaches. By employing a Transformer-based architecture, our method efficiently processes motion and text embeddings, enabling the generation of motion sequences that exhibit strong alignment with the provided textual descriptions.
In this section, we provide a concise overview of flow matching, followed by the details of our proposed CFM-based framework, including the training procedure and sampling strategy.
3.1 Flow Matching
The flow matching [37] determines a time-dependent vector field $v:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, that transforms a simpler probability density function $p_{0}$, such as a Gaussian distribution, into a more complex one, $p_{1}$, through the so-called probability density path $p:~{}[0,1]~{}\times~{}\mathbb{R}^{d}~{}\rightarrow~{}\mathbb{R}^{+}$, where $p_{0}=p(0,\cdot)$ at the initial time $t=0$, and $p_{1}=p(1,\cdot)$ at the final time $t=1$.
The vector field $v$ defines the ordinary differential equation (ODE):
$\frac{d}{dt}\phi(t,x)$ $=v(t,\phi(t,x)),$ (1)
$\phi(0,x)$ $=x,$
where its solution $\phi:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is named flow, the diffeomorphism induced by the vector field $v$. This ensures that $\phi$ possesses a differentiable inverse, guaranteeing a smooth and invertible mapping of the probability space. For notational convenience, we use $\phi_{t}$ to denote $\phi(t,\cdot)$ and $v_{t}$ to denote $v(t,\cdot)$.
According to Chen et al. [57], one can reparameterize the vector field $v$ using a neural network with parameters $\theta\in\mathbb{R}^{\ell}$. Consequently, the flow $\phi$ is also parameterized by $\theta$, resulting in a Continuous Normalizing Flow (CNF). This CNF transforms the initial density $p_{0}$ to the density $p_{t}$ at time $t$ via a push-forward operation, specifically a change of variables:
$p_{t}=[\phi_{t}]_{*}p_{0}.$
This transformation is defined as:
$p_{t}(y)=p_{0}(x)\left|\det\left(\frac{\partial\phi_{t}^{-1}(y)}{\partial y}\right)\right|,$
where $x=\phi_{t}^{-1}(y)$ for $t\in[0,1]$.
Given a finite set of samples $x_{1}$ from an unknown data distribution $q$, and initializing with $p_{0}=\mathcal{N}(0,I)$, Lipman et al. [37] introduce the concept of a conditional probability path $p_{t}(\cdot|x_{1}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$, defined for each sample $x_{1}$. At the final time $t=1$, $p_{1}(\cdot|x_{1})$ is defined as a Gaussian distribution centered at $x_{1}$ with a small standard deviation $\sigma_{\min}\geq 0$, concentrating the probability around the sample. The distribution $p_{t}(x)$ is then defined as the marginalization of these conditional probability paths:
[…]
Thus, at $t=1$, the marginal distribution $p_{1}$ approximates the unknown distribution $q$.
However, the direct computation of $p_{t}(x)$ via Eq. (2) is intractable. Lipman et al. [37] demonstrate that an objective function designed to approximate the unknown vector field $v_{t}$, which generates the marginal probability path $p_{t}$, has gradients identical, with respect to the model parameters, to an objective function that approximates the conditional vector field $u_{t}(\cdot|x_{1})$. This crucial property avoids the intractability of computing $p_{t}(x)$. Consequently, it suffices to define appropriate conditional probability paths $p_{t}(\cdot|x_{1})$ and conditional vector fields $u_{t}(\cdot|x_{1})$ to minimize the objective function with respect to these conditional vector fields.
Thus, for each $x_{1}\sim q$, a conditional Gaussian probability path is defined for each time $t\in[0,1]$:
$p_{t}(x|x_{1})=\mathcal{N}(x;\mu_{t}(x_{1}),\sigma_{t}(x_{1})^{2}I),$
where, as defined previously, at time $t=1$: $p_{1}(x|x_{1})=\mathcal{N}(x;x_{1},\sigma_{\text{min}}^{2}I)$, with $\mu_{1}(x_{1})=x_{1}$ and $\sigma_{1}(x_{1})=\sigma_{\text{min}}$. Similarly, at $t=0$, $p_{0}(x|x_{1})=\mathcal{N}(x;0,I)$, where $\mu_{0}(x_{1})=0$ and $\sigma_{0}(x_{1})=~{}1$, corresponding to the standard normal distribution. Finally, for $10<t<$ 1, we define a conditional probability path where the mean and standard deviation are linearly interpolated between the boundary parameters $\{\mu_{0},\sigma_{0}\}$ and $\{\mu_{1},\sigma_{1}\}$:
$\mu_{t}(x_{1})$ $=tx_{1},$
$\sigma_{t}(x_{1})$ $=1-(1-\sigma_{\text{min}})t.$
This results in the general form:
$p_{t}(x|x_{1})=\mathcal{N}\left(x;tx_{1},\left(1-(1-\sigma_{\text{min}})t\right)^{2}I\right).$
Under the Gaussian probability path $p_{t}$, an element from the range of the flow at time $t$ is given by $\psi_{t}(x)\in\mathbb{R}^{d}$, which is a function of the linear parameters $\mu_{t}(x_{1})$ and $\sigma_{t}(x_{1})$, as follows:
$\psi_{t}(x)=\left(1-(1-\sigma_{min})t\right)x+tx_{1}.$ (3)
The original Conditional Flow Matching (CFM) objective [37] is formulated with respect to the initial sample $x_{0}$. However, for our iterative sampling process, we propose redefining the objective in terms of the intermediate state $x_{t}$. Utilizing the conditional vector field construction from Lipman et al. [37], which expresses the vector field based on the means, standard deviations, and their derivatives. Thus, we adopt their formulation to define the conditional vector field:
$u_{t}(x_{t}|x_{1})=\frac{x_{1}-(1-\sigma_{min})x_{t}}{1-(1-\sigma_{min})t},$ (4)
where $x_{t}=\psi_{t}(x)$.
Rather than training a model to approximate $u_{t}$ directly, we propose predicting the target $x_{1}$ through a neural network conditioned on $x_{t}$ (Sec. 3.2.1). Given the predicted $x_{1}$ and the fixed $\sigma_{\text{min}}$, the vector field $u_{t}$ is computed via Eq. (4). This approach bypasses explicit vector field estimation while preserving the theoretical guarantees of CFM.
3.2 Framework
We start from the conditional probability path proposed by Lipman et al. [37], as detailed previously, along with its associated flow and conditional vector field. These constructs are based on the interpolation of parameters of Gaussian distributions along a temporal trajectory defined by $t\in[0,1]$.
Crucially, these conditional formulations depend on the samples $x_{1}$, which in our case correspond to human motion sequences extracted from our datasets. Specifically, $x_{1}\in\mathbb{R}^{N\times J\times D}$, where $N$ represents the number of poses in the sequences, $J$ the number of joints per pose, and $D$ the dimensionality of the features for each joint. This entire framework operates under a condition $c$, which, in our context, corresponds to the encoded text (Fig. 2).
While our model implements the architecture of MDM [20], it is crucial to highlight that the input $x_{t}$ is derived from a linear interpolation process, following the CFM framework (Eq. (3)), rather than the stochastic corruption process characteristic of diffusion models such as MDM.
Figure 2: Text-Driven Motion Generation: Leveraging the architecture proposed by Tevet et al. [20], our approach generates motion sequences from an input ${x}_{t}=({x}^{1}_{t},{x}^{2}_{t},\dots,{x}^{N}_{t})$, where each ${x}^{i}_{t}\in\mathbb{R}^{J\times D}$ denotes the pose of the $i$-th frame. Crucially, this input is derived via conditional flow matching. By processing this input through a Transformer Encoder, the model produces a motion sequence ${x}_{1}=({x}^{1}_{1},{x}^{2}_{1},\dots,{x}^{N}_{1})$.
The proposed framework allows us to define a training objective that directly targets motion, thereby mitigating jittering, as outlined in the following.
3.2.1 Training
The training process begins by taking a sample $x_{1}$ from our dataset and simultaneously sampling $x_{0}$ from the initial Gaussian distribution. We then utilize the flow defined in Eq. (3) to transport $x_{0}$ along a trajectory conditioned by $x_{1}$, enabling us to compute the corresponding point $x_{t}$ at time $t\in[0,1]$. Specifically:
$x_{t}=\psi_{t}(x_{0})=(1-(1-\sigma_{min})t)x_{0}+tx_{1},$
where $x_{0}\sim\mathcal{N}(0,I)$ (Fig. 3).
Following a similar approach proposed by Ramesh et al. [49] for diffusion models, where they modified the loss function to predict the target $x_{1}$ directly—achieving improved performance compared to predicting the noise—we tailor this concept to our CFM framework. Thus, our training objective is defined as:
$\mathbb{E}_{x_{1}\sim q(x_{1}|c),t\sim\mathcal{U}[0,1]}\lVert\mathcal{G}_{p}(x_{t},t,c;\theta)-x_{1}\lVert^{2}_{2},$ (5)
where $\mathcal{U}[0,1]$ denotes a uniform distribution, $\mathcal{G}_{p}$ is our trainable model, and $\theta$ represents its learnable parameters.
Empirically, this formulation reduces motion jitter and improves generalization compared to prior CFM and diffusion-based methods, as validated in Sec. 5.
Figure 3: Overview of the training process. In each epoch, the process starts with a sample $x_{1}$ from the training dataset and a sample $x_{0}\sim\mathcal{N}(0,{I})$. An intermediate representation $x_{t}$ is then determined via linear interpolation between $x_{1}$ and $x_{0}$. The region highlighted in red denotes the space of valid human motions. Note that at the beginning of the process, both $x_{0}$ and $x_{t}$ can be found outside this space.
3.3 Model
Our model leverages a Transformer Encoder for motion processing, a technique proven effective for human motion synthesis [41, 58, 59]. Adopting the architecture of Tevet et al. [20], our approach modifies the source of the input $x_{t}$. Specifically, we employ flow matching to generate $x_{t}$, while maintaining three key inputs:
  • 1.
    Textual descriptions encoded via a frozen CLIP model.
  • 2.
    Time embeddings through learned positional encoding.
  • 3.
    The motion sequence $x_{t}$, projected to the latent dimension of the Transformer.
Text and time encodings are first summed in the latent space, then concatenated with $x_{t}$ to form the combined input representation (Fig. 2). This fused input is processed by the Transformer to predict the target motion $x_{1}$ via the objective function (Eq. (5)), maintaining the temporal resolution of the input while removing noise artifacts. The end-to-end design enables precise control over motion smoothness while preserving fidelity to text conditions.
Table 1: Comparison of methods on the HumanML3D dataset. Red highlights the best values, and blue indicates the second-best values. Arrows indicate the desired direction: $\uparrow$ higher is better, $\downarrow$ lower is better, $\rightarrow$ closer to the target value is better.
Methods RP Top3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MMModality $\uparrow$ Jitter $\rightarrow$
Real motion $0.797^{\pm.001}$ $0.001^{\pm.000}$ $2.973^{\pm.005}$ $9.488^{\pm.081}$ $47.262$^{\pm$0.069$}$
TM2T [43] $0.726^{\pm.002}$ $1.507^{\pm.019}$ $3.469^{\pm.011}$ $8.526^{\pm.086}$ $2.540^{\pm.063}$ $87.208$^{\pm$0.151$}$
T2M-GPT [21] $0.776^{\pm.005}$ $0.124^{\pm.008}$ $3.131^{\pm.019}$ $9.567^{\pm.101}$ $1.719^{\pm.120}$ $76.046$^{\pm$1.123$}$
MDM [20] $0.706^{\pm.004}$ $0.519^{\pm.050}$ $3.658^{\pm.022}$ $9.442^{\pm.070}$ $2.871^{\pm.047}$ $46.30^{\pm.9}$
MotionGPT [60] $0.659^{\pm.003}$ $0.185^{\pm.007}$ $4.013^{\pm.022}$ $9.291^{\pm.067}$ $3.481^{\pm.119}$ $94.770$^{\pm$0.250$}$
MotionLCM [35] $0.802^{\pm.002}$ $0.416^{\pm.010}$ $3.008^{\pm.005}$ $9.740^{\pm.069}$ $2.123^{\pm.076}$ $37.578$^{\pm$0.224$}$
MLD [36] $0.758^{\pm.003}$ $0.427^{\pm.012}$ $3.269^{\pm.015}$ $9.775^{\pm.073}$ $2.573^{\pm.089}$ $22.772$^{\pm$0.113$}$
MFM [39] $0.666^{\pm.005}$ $0.446^{\pm.047}$ $5.186^{\pm.023}$ $9.987^{\pm.089}$ $2.419^{\pm.096}$ $576.679$^{\pm$3.153$}$
FlowMotion-Big (ours) $0.648^{\pm.004}$ $0.268^{\pm.028}$ $5.319^{\pm.032}$ $9.796^{\pm.069}$ $2.327^{\pm.077}$ $40.81^{\pm.5}$
FlowMotion (ours) $0.663^{\pm.005}$ $0.278^{\pm.030}$ $5.239^{\pm.027}$ $9.788^{\pm.103}$ $2.349^{\pm.055}$ $39.522$^{\pm$0.770$}$
3.4 Sampling
To generate the final point $x_{1}$ at time $t=1$, the sampling process begins with $x_{0}\sim\mathcal{N}(0,I)$ at time $t=0$. We then propagate $x_{0}$ through the flow $\psi$, as defined in Sec. 3.2.1. This involves solving Eq. (1) with initial condition $\psi_{0}(x_{0})=x_{0}$ to obtain $\psi_{1}(x_{0})$. We approximate the solution via Euler integration [39], yielding the iterative update:
$x_{i+1}=x_{i}-h\left(u_{t_{i}}(x_{i}|x_{1})\right),$ (6)
where $i\in\{0,1,\dots,M-1\}$, $M$ denotes the number of iterations, $h=1/M$ represents the step size, and the time instance is given by $t_{i}=i/M$. Observe that $t_{i}\in[0,1)$. The conditional vector field $u_{t_{i}}$, defined in Eq. (4), depends on $x_{1}$ and $x_{i}$.
To enhance the influence of the textual condition $c$, we introduce a guided model $\mathcal{G}$ based on the classifier-free guidance technique. Leveraging our trained model $\mathcal{G}_{p}$, which predicts $x_{1}$, the guided model $\mathcal{G}$ is defined as:
$\mathcal{G}(x_{t},t,c;\theta)=\mathcal{G}_{p}(x_{t},t,\emptyset;\theta)+s\left(\mathcal{G}_{p}(x_{t},t,c;\theta)-\mathcal{G}_{p}(x_{t},t,\emptyset;\theta)\right),$
where $s$ is the guidance scale. This formulation interpolates between conditional (with text input $c$) and unconditional (where $c=\emptyset$) generation, allowing us to control the influence of the textual condition.
Thus, substituting $\mathcal{G}$ into Eq. (4) and replacing into Eq. (6), we obtain the iterative sampling process:
$x_{i+1}=x_{i}-h\Bigg{[}\frac{\mathcal{G}(x_{i},t_{i},c;\theta)-(1-\sigma_{\text{min}})x_{i}}{1-(1-\sigma_{\text{min}})t_{i}}\Bigg{]}.$
As previously mentioned, the objective of flow matching, and in our specific case CFM, is to determine the conditional vector field $u_{t}$. This is achieved from the model designed to estimate the target directly. By obtaining $\mathcal{G}$ and having the previous iterative value $x_{i}$, we can derive the desired vector field using Eq. (4) without the need to generate the vector field directly.
3.5 Implementation Details
The Transformer Encoder architecture utilizes a hidden dimension of 512. The text encoder comprises 8 layers, 4 attention heads, and a feedforward hidden layer dimension of 1024. We employ the pretrained CLIP-ViT-B/32 model for text processing, keeping its parameters frozen throughout training. Our model was trained with a batch size of 128.
During sampling, we set $\sigma_{min}$ to zero. We use Euler integration with 100 steps and a classifier-free guidance scale of 2.5, a standard value in similar works [20]. For optimization, we adopted the AdamW [61] optimizer, configured with parameters $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999, and a learning rate set to $1\times 10^{-4}$.
4 Experiments
All experiments were conducted on a Linux system equipped with an NVIDIA A40 GPU card. The training process typically took about 20 hours to complete, whereas generating samples required approximately 2 seconds per instance. This represents a significant speed advantage compared to diffusion models as MDM [20], which typically require about 30 seconds for sampling, and also surpasses other existing diffusion-based models in terms of sampling speed [19].
Our model generates human motion sequences conditioned on textual descriptions. To quantitatively assess the quality and diversity of the generated motions, we adopt a range of metrics proposed by Guo et al. [30] for evaluating text-conditioned human motion generation. These metrics allow us to compare our approach to state-of-the-art methods across various dimensions, including the fidelity of the motion to the text, the diversity of the generated motions, and the smoothness of the motion.
4.1 Datasets
We evaluate FlowMotion on two benchmark datasets: HumanML3D [30] and KIT [44].
  • 1.
    HumanML3D: This dataset, introduced by Guo et al. [30], combines the HumanAct12 [62] and AMASS [63] datasets. It consists of 14,616 motion sequences at 20 frames per second. The original motion sequences can exceed 10 seconds in duration, but following prior work [30], they are randomly truncated to a maximum of 10 seconds (196 frames) for consistency, with a minimum sequence length of 40 frames. The dataset is paired with 44,970 textual descriptions. HumanML3D is generally considered the primary benchmark for assessing the performance of different approaches in this task. The results of our experiments on this dataset are detailed in Table 1.
  • 2.
    KIT Motion-Language Dataset [44]: This dataset provides a valuable resource for evaluating text-conditioned motion generation, containing 3,911 motion sequences and 6,353 textual descriptions that capture a diverse range of human actions and movements. Motion sequences have a maximum length of 196 frames and a minimum of 24 frames. The results for this dataset are presented in Table 2.
Table 2: Comparison of methods on the KIT dataset. Red highlights the best values, and blue indicates the second-best values. Arrows indicate the desired direction: $\uparrow$ higher is better, $\downarrow$ lower is better, $\rightarrow$ closer to the target value is better.
Methods RP Top3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MMModality $\uparrow$ Jitter $\rightarrow$
Real motion $0.784^{\pm.003}$ $0.026^{\pm.003}$ $2.772^{\pm.013}$ $11.016^{\pm.095}$ $49.922$^{\pm$0.165$}$
T2M-GPT [21] $0.737^{\pm.004}$ $0.469^{\pm.010}$ $3.002^{\pm.013}$ $11.006^{\pm.111}$ $1.903^{\pm.070}$ $98.425$^{\pm$0.291$}$
MDM [20] $0.731^{\pm.004}$ $0.505^{\pm.027}$ $3.077^{\pm.017}$ $10.705^{\pm.098}$ $1.782^{\pm.152}$ $73.21^{\pm.5}$
MFM [39] $0.405^{\pm.004}$ $0.327^{\pm.016}$ $9.155^{\pm.028}$ $10.707^{\pm.070}$ $1.639^{\pm.137}$ $1099.988$^{\pm$1.183$}$
FlowMotion (ours) $0.404^{\pm.005}$ $0.396^{\pm.042}$ $9.206^{\pm.022}$ $10.989^{\pm.082}$ $1.756^{\pm.083}$ $52.40^{\pm.3}$
4.2 Motion Representation
Consistent with Guo et al. [30], we represent each pose using the features $x_{f}=(\dot{r}_{a},\dot{r}_{x},\dot{r}_{z},r_{y},j_{p},j_{v},j_{r},c_{f})$, where $\dot{r}_{a}\in\mathbb{R}$ denotes the global root angular velocity, $\dot{r}_{x},\dot{r}_{z}\in\mathbb{R}$ represent the global root velocity in the X-Z plane, $r_{y}$ is the root height, $j_{p}\in\mathbb{R}^{3j}$, $j_{v}\in\mathbb{R}^{3j}$, $j_{r}\in\mathbb{R}^{6j}$ correspond to the local joint positions, velocities, and rotations respectively, with $j$ being the number of joints, and $c_{f}\in\mathbb{R}^{4}$ represents the foot contact features, derived from the heel and toe joint velocities. Thus, each frame within our motion sequences is characterized by this set of features, $x_{f}$, meaning that each $x_{t}$ input to our model comprises frames with this format.
These motion features encapsulate local joint positions, velocities, and rotations within the root space, in addition to global translation and rotation. The dimensionality of these features is directly determined by the number of joints considered. Specifically, for the HumanML3D dataset, which uses $j=22$ joints, the resulting feature dimension is 263. The KIT-ML dataset, with $j=21$, has a feature dimension of 251. This difference in dimensionality reflects the varying complexity captured by the motion representation of each dataset.
4.3 Metrics
The evaluation metrics employed are those proposed by Guo et al. [30], which have become standard practice in the human motion generation literature [21, 19, 55, 39, 20]. According to established evaluation protocols, such as those used by Guo et al. [30], each metric is computed over 20 trials to account for the inherent stochasticity in the sampling and evaluation process, and we report the average and standard deviation across these trials, as shown in Tables 12. The metrics used to evaluate 3D human motion generation are as follows:
Fréchet Inception Distance (FID)
Quantifies the similarity between the distribution of generated motions and the distribution of ground-truth motions from the test set, where both sets of motions are conditioned on the same textual descriptions from the test set.
R-Precision (RP Top-k)
Measures the correspondence between the generated motion and the input text. We report Top-1, Top-2, and Top-3 R-Precision, representing the probability of retrieving the correct motion given the text within the top 1, 2, or 3 generated samples, respectively.
Multimodal Distance (MM-Dist)
Calculates the average distance between a set of randomly sampled text embeddings and their corresponding generated motion embeddings.
Diversity
Assesses the variability of the generated motions. Motions are encoded into a shared latent space, and the average pairwise distance between randomly sampled motions is calculated. A greater average distance signifies higher diversity.
Multimodality (MModality)
Determines the average Euclidean distance between generated motions encoded in the same latent space when conditioned on the identical text.
Jitter
Evaluates motion smoothness by measuring the jerk (rate of change of the acceleration) of each joint. Following Du et al. [4], we compute it as the average magnitude of the acceleration of body joints. Lower jitter values indicate smoother motion. To ensure comparability across datasets, we introduce a scaling factor, denoted as $\alpha$, derived from the ratio of motion ranges of the HumanML3D and KIT datasets. The motion range for each dataset is calculated as the difference between the mean of the maximum joint positions and the mean of the minimum joint positions across all motion sequences. Formally, the scaling factor $\alpha$ is computed as:
$\alpha=\frac{\text{range}_{\text{humanml3d}}}{\text{range}_{\text{kit}}}=\frac{\bar{x}^{\text{max}}_{\text{humanml3d}}-\bar{x}^{\text{min}}_{\text{humanml3d}}}{\bar{x}^{\text{max}}_{\text{kit}}-\bar{x}^{\text{min}}_{\text{kit}}},$
where $\bar{x}^{\text{max}}_{d}$ and $\bar{x}^{\text{min}}_{d}$ represent the mean of the maximum and minimum joint positions for dataset $d$, respectively. Empirically, $\alpha$ was determined to be 0.00073. This scaling factor is applied to modulate the magnitude of jitter values obtained from the KIT dataset using the formula:
$\text{Jitter}_{\text{kit}}^{\text{scaled}}=\alpha\cdot\text{Jitter}_{\text{kit}}.$
This normalization mitigates the tendency for jitter values to become excessively large on the KIT dataset, ensuring a consistent perturbation magnitude across datasets and enhancing robustness and comparability.
To facilitate the comparison of textual and motion data, both are encoded into a shared latent space. This is enabled by the pretrained motion and text encoders provided by Guo et al. [30], allowing for quantitative evaluation based on Euclidean distances in this embedding space.
4.4 Model Parameter Analysis
Our model is based on a Transformer Encoder, as introduced by Vaswani et al. [64]. These networks allow for flexibility in adjusting parameters like the feedforward dimension and the number of attention heads. Drawing inspiration from the original Transformer work, a scaled-up version is also explored. We introduce two variants:
  • 1.
    FlowMotion (Standard): 8 layers, 4 attention heads, 1024 feedforward dimension
  • 2.
    FlowMotion-Big: 8 layers, 16 attention heads, 2048 feedforward dimension
The scaled-up FlowMotion-Big variant improves the FID-Jitter balance marginally (Table 1); however, this comes at the cost of increased model complexity and computational overhead. Both variants retain the original Transformer 512-dimensional hidden representation in the encoder, consistent with prior work [64].
Figure 4: Comparison of motion trajectories generated by MDM [20], MFM [39], and the FlowMotion Model. MDM and MFM fail to depict a complete circular trajectory accurately. In contrast, the FlowMotion Model successfully generates a motion sequence in which the starting and ending points coincide. This result demonstrates the superior ability of the FlowMotion Model to interpret the motion instruction and generate a complete circular trajectory.
5 Results
Our model achieves optimal balance between FID and Jitter, demonstrating exceptional Jitter performance that approximates lifelike human motion. It outperforms all other methods in this regard, except for the diffusion model MDM, which exhibits a slightly better Jitter score but suffers from a higher FID and a significantly slower sampling speed (Tables 12). This balance suggests that our model is especially suitable for applications requiring both performance and high-quality motion generation.
On the HumanML3D dataset, our model achieves better FID than both the diffusion-based method MDM and the flow matching method MFM. While some models, such as MLD [36] and MotionLCM [60], achieve lower Jitter scores than ours, they achieve this with a substantial compromise in generalization capability, reflected in higher FID values (0.41 compared to our 0.27, Table 1). The lower Jitter scores, surpassing those of the ground truth, can potentially be attributed to factors such as differences in the generated sequence length or interpolation strategies, which may result in artificially smoother transitions. Conversely, models that exhibit better generalization than ours, such as the Vector Quantized Variational Autoencoder (VQ-VAE) based T2M-GPT and MotionGPT, tend to produce less smooth motion sequences, with Jitter values approximately double that of the ground truth (Tables 1-2).
Notably, our model maintains a low FID, indicative of strong generalization, and remarkably low Jitter scores on the HumanML3D dataset. If the goal is to generate motions as smooth as those in the ground truth data, our model achieves the second-best Jitter score, with MDM being the only model with a slightly lower Jitter (46 compared to the ground truth value of 47). However, MDM exhibits a much higher FID (0.52 compared to our 0.27), indicating a deficiency in generalization that our model successfully addresses (Fig. 4).
On the KIT dataset, our method achieves FID values generally around 0.4, coupled with the best Jitter score compared to all methods (Table 2), highlighting its ability to generate smooth, high-quality motion sequences while maintaining excellent generalization capability across different datasets. This superior balance is also visually apparent when comparing the generated motions, as shown in Fig. 5, which depicts a motion sequence from a side profile. As shown in Fig. 5, MFM produces motions with noticeable jerkiness. Our method, while having the same number of frames as MFM, exhibits smoother transitions throughout the motion, making it appear more natural. While matching MFM in the number of frames, our approach yields significantly more natural transitions, underscoring its effectiveness in generating high-fidelity motion.
Figure 5: Qualitative comparison of motion sequences generated from the text prompt “a person walks forward with wide steps,” comparing MFM [39], T2M-GPT [21], and FlowMotion. The capture shows the sequence of movements taken from the side profile, with foot traces visible below.
Further evidence of robustness of our model, when trained on KIT-ML [44], is presented in Figure 6, directly comparing motion generation against MFM [39]. A key advantage, enhanced motion stability, is visually highlighted: Figure 6a shows MFM generation with noticeable tremors and motion artifacts, including body part instability, even during intended stillness before a step. In contrast, captured at the same instant (Figure 6b), our FlowMotion model showcases stability, maintaining a coherent pose without involuntary movements or tremors. This visual comparison underscores improved stability and reduced artifact generation achieved by FlowMotion.
Figure 6: Motion generation comparison for KIT-ML trained models with prompt “a person walks forwards and stops” at pre-step instant: (a) MFM: tremors and motion artifacts including body part distortions, visible within green demarcations (b) FlowMotion: stable and coherent pose
We also analyze the trade-off between motion quality and sampling steps, as illustrated in Fig. 7, which shows that both FID and Jitter values exhibit a trend towards stabilization after approximately 50 sampling steps, approaching the error region of our previously reported results (horizontal lines in Fig. 7). A more complete set of metrics, including diversity and R-precision, is presented in Table 3 for sampling steps ranging from 1 to 100. Notably, all metrics remain relatively stable from around 50 steps onwards, consistent with our results in Table 1. This analysis, conducted on the HumanML3D test set, indicates that a sampling regime of around 50 steps is sufficient for generating high-quality motions with good generalization, striking a balance between computational cost and motion fidelity.
Figure 7: Comparison of FID and Jitter for our model across 100 sampling steps, taken every 5 steps. The graph displays results from the 5th step onwards for better visualization. The horizontal blue band represents the FID value of 0.278 $\pm$ 0.03, and the horizontal red band represents the Jitter value of 39.5 $\pm$ 0.77, both obtained by our model as reported in Table 1.
Further analysis of our model’s performance under various guidance strengths reveals an optimal balance at a guidance scale of 2.5 (Fig. 8). At this setting, the model maintains an FID score below 0.3 while achieving a Jitter score near 40, approaching the ground truth value of 47. This configuration, which we employed for our model’s generation, demonstrates the effectiveness of carefully tuning the guidance scale to achieve both high fidelity and smooth motion quality.
Figure 8: FID and Jitter metrics across varying guidance scales for our model, ranging from 0.5 to 5 in increments of 0.5.
Moreover, our model demonstrates a more precise interpretation of textual prompts when compared to MDM and MFM (Fig. 4). Notably, it successfully generates a complete circular trajectory with coinciding start and end points, a feat not achieved by MDM or MFM.
Table 3: Comparison of metrics for different numbers of sampling steps (NS) on the HumanML3D dataset. Red highlights the best values, and blue indicates the second-best values. Arrows indicate the desired direction: $\uparrow$ higher is better, $\downarrow$ lower is better, $\rightarrow$ closer to the target value is better.
NS R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ Jitter $\rightarrow$
Top-1 Top-2 Top-3
1 $0.294^{\pm.005}$ $0.463^{\pm.005}$ $0.581^{\pm.004}$ $2.515^{\pm.065}$ $5.601^{\pm.023}$ $8.992^{\pm.075}$ $188.013$^{\pm$1.751$}$
5 $0.318^{\pm.005}$ $0.492^{\pm.005}$ $0.607^{\pm.005}$ $1.528^{\pm.071}$ $5.622^{\pm.026}$ $9.485^{\pm.080}$ $47.40^{\pm 0.3}$
10 $0.353^{\pm.006}$ $0.535^{\pm.009}$ $0.645^{\pm.007}$ $0.610^{\pm.033}$ $5.360^{\pm.032}$ $9.757^{\pm.094}$ $38.567$^{\pm$0.300$}$
20 $0.375^{\pm.006}$ $0.557^{\pm.007}$ $0.663^{\pm.006}$ $0.413^{\pm.040}$ $5.239^{\pm.033}$ $9.926^{\pm.100}$ $38.074$^{\pm$0.506$}$
30 $0.373^{\pm.005}$ $0.558^{\pm.006}$ $0.667^{\pm.005}$ $0.371^{\pm.037}$ $5.237^{\pm.029}$ $9.807^{\pm.086}$ $36.310$^{\pm$0.611$}$
40 $0.373^{\pm.006}$ $0.553^{\pm.007}$ $0.654^{\pm.006}$ $0.324^{\pm.041}$ $5.253^{\pm.033}$ $9.778^{\pm.092}$ $38.488$^{\pm$0.498$}$
50 $0.372^{\pm.004}$ $0.554^{\pm.007}$ $0.659^{\pm.006}$ $0.295^{\pm.039}$ $5.263^{\pm.037}$ $9.803^{\pm.089}$ $39.190$^{\pm$0.543$}$
60 $0.373^{\pm.007}$ $0.554^{\pm.006}$ $0.664^{\pm.007}$ $0.289^{\pm.028}$ $5.234^{\pm.037}$ $9.906^{\pm.072}$ $39.056$^{\pm$0.504$}$
70 $0.369^{\pm.007}$ $0.555^{\pm.006}$ $0.663^{\pm.006}$ $0.293^{\pm.032}$ $5.236^{\pm.025}$ $9.796^{\pm.080}$ $39.084$^{\pm$0.541$}$
80 $0.377^{\pm.005}$ $0.558^{\pm.005}$ $0.664^{\pm.006}$ $0.299^{\pm.032}$ $5.234^{\pm.021}$ $9.770^{\pm.077}$ $39.021$^{\pm$0.641$}$
90 $0.370^{\pm.004}$ $0.551^{\pm.005}$ $0.658^{\pm.005}$ $0.277^{\pm.029}$ $5.262^{\pm.030}$ $9.763^{\pm.067}$ $38.792$^{\pm$0.615$}$
100 $0.370^{\pm.006}$ $0.553^{\pm.007}$ $0.663^{\pm.005}$ $0.278^{\pm.030}$ $5.239^{\pm.027}$ $9.788^{\pm.103}$ $39.52^{\pm 0.7}$
Real Motion $0.511^{\pm.003}$ $0.703^{\pm.002}$ $0.796^{\pm.002}$ $0.001^{\pm.000}$ $2.972^{\pm.008}$ $9.481^{\pm.107}$ $47.26^{\pm 0.1}$
6 Conclusions
We introduce FlowMotion, a text-to-motion generation framework leveraging Conditional Flow Matching (CFM). Our key contribution is a novel training objective that directly predicts target motions, bypassing intermediate velocity estimation to reduce jitter and improve stability. Evaluations on the HumanML3D and KIT datasets demonstrate that FlowMotion achieves state-of-the-art fidelity (FID) and smoothness (jitter) while maintaining competitive efficiency against diffusion-based and flow matching baselines. The sampling process generates motion sequences in approximately 2 seconds—significantly faster than traditional diffusion models—enabling practical deployment in resource-constrained scenarios.
FlowMotion balances computational efficiency and motion quality without relying on oversized architectures. While some methods achieve marginally lower jitter, our framework prioritizes a holistic trade-off between perceptual realism, temporal consistency, and speed. Future work will explore enforcing physical plausibility (e.g., ground contact constraints to eliminate floating artifacts) and extending the framework to tasks like motion editing and full-body articulation. The proposed training objective also opens avenues for broader applications in flow-matching frameworks beyond human motion synthesis.
Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) - Finance Code 001.
Declaration of Generative AI in Scientific Writing
In the process of preparing this manuscript, the authors employed ChatGPT-4 and Gemini 2.0 to refine the language and enhance the overall readability. After using this tools, the authors meticulously reviewed and revised the content as necessary and assume full responsibility for the final content of this publication.